# Asynchronous Predictive Counterfactual Regret Minimization+ Algorithm in Solving Extensive-Form Games

### v1

@@ Appendix B An Alternative Regret Upper Bound of APCFR+
$\footnotesize R^{T}(I)\leq\sqrt{\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}}.$
@@
$\footnotesize R^{T}(I)\leq\min\left(\sqrt{{\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)}},\ \sqrt{\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}}\right).$
We employ Theorem 4.1 rather than Theorem B.1 in the main text because the regret upper bound presented in Theorem B.1 is typically significantly larger than that in Theorem 4.1. In our experiments (as shown in Figures 12 and 13), we observe that the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ consistently increases over time. In contrast, the value of $\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)$ tends to stabilize, exhibiting a flattening trend. We hypothesize that this phenomenon arises from the fact that even when ${\bm{r}}^{t}_{I}$ and $\bm{r}^{t-1}_{I}$ are close, the value of $\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ remains extremely large.
@@
$\leq$ $\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+\frac{1}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
@@
$\leq$ $\langle{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\frac{1}{1+\alpha^{t}_{I}}\langle\bm{r}^{t-1}_{I},\hat{\bm{\theta}}^{t}_{I}-{\bm{\theta}}^{t+1}_{I}\rangle+\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})\right)+\frac{1}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right).$
Combining Eq. (26) and (35), we have
$R^{T}(I)\leq$ $\frac{1}{\eta}\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})+\sum^{T}_{t=1}\langle{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\sum^{T}_{t=1}\frac{1}{1+\alpha^{t}_{I}}\langle\bm{r}^{t-1}_{I},\hat{\bm{\theta}}^{t}_{I}-{\bm{\theta}}^{t+1}_{I}\rangle+\sum^{T}_{t=1}\frac{1}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$ (36)
@@
where the third inequality comes from that $\forall a,b,\rho>0,ab\leq a^{2}/(2\rho)+\rho b^{2}/2$ ($a=\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|_{2}$, $b=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}$, and $\rho=1/\eta$), and the last inequality is from $\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|^{2}_{2}/2$. Then, from Eq. (29) ($\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})\leq\frac{1}{2}$), we get
@@
Appendix C Details of APDCFR+ and Stable APDCFR+
We now detail APDCFR+ and Stable APDCFR+. Before presenting these algorithms, we first introduce PDCFR+ (Xu et al., 2024b). At each iteration $t$, PDCFR+ updates the strategy at each infoset $I\in\mathcal{I}$ according to
$\hat{\bm{R}}^{t}_{I}=\left[\frac{t^{\beta}}{1+t^{\beta}}\bm{R}^{t}_{I}+\bm{r}^{t-1}_{I}\right]^{+},\quad\bm{R}^{t+1}_{I}=\left[\frac{t^{\beta}}{1+t^{\beta}}\bm{R}^{t}_{I}+\bm{r}^{t}_{I}\right]^{+},$ (40)
$\sigma^{t}_{i}(I)=\frac{\left[\hat{\bm{R}}^{t}_{I}\right]^{+}}{\|\left[\hat{\bm{R}}^{t}_{I}\right]^{+}\|_{1}}=\frac{\hat{\bm{R}}^{t}_{I}}{\|\hat{\bm{R}}^{t}_{I}\|_{1}},$
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, $\bm{r}^{0}_{I}=\bm{0}$, and $\beta=2.3$ (as presented in the Experiments section of Xu et al. (2024b)).
Next, we introduce APDCFR+, which combines APCFR+ and PDCFR+. As noted in the main text, APDCFR+ employs a novel aggressive discounting scheme for accumulated regrets, in contrast to the original PDCFR+. At each iteration $t$, the strategy is updated at each infoset $I\in\mathcal{I}$ as follows:
$\hat{\bm{R}}^{t}_{I}=\left[\frac{t^{\beta}}{\kappa+t^{\beta}}\bm{R}^{t}_{I}+\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\right]^{+},\ \bm{R}^{t+1}_{I}=\left[\frac{t^{\beta}}{\kappa+t^{\beta}}\bm{R}^{t}_{I}+\bm{r}^{t}_{I}\right]^{+},$ (41)
$\sigma^{t}_{i}(I)=\frac{\left[\hat{\bm{R}}^{t}_{I}\right]^{+}}{\|\left[\hat{\bm{R}}^{t}_{I}\right]^{+}\|_{1}}=\frac{\hat{\bm{R}}^{t}_{I}}{\|\hat{\bm{R}}^{t}_{I}\|_{1}},\ \alpha^{t}_{I}=\min\left(\sqrt{\frac{\sum_{\tau=1}^{t-1}\|\bm{r}^{\tau}_{I}-\bm{r}^{\tau-1}_{I}\|^{2}_{2}}{\sum_{\tau=1}^{t-1}\|\bm{R}^{\tau+1}_{I}-\bm{R}^{\tau}_{I}\|^{2}_{2}}},10\right),$
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, and $\bm{r}^{0}_{I}=\bm{0}$. Note that in the last line, we use $10$ instead of $5$, as we observe that $10$ performs better in HUNL Subgames. In addition, we apply the term $\frac{t^{\beta}}{\kappa+t^{\beta}}$ instead of $\frac{t^{\beta}}{1+t^{\beta}}$ for discounting the accumulated regrets, where $\kappa=500$ is set to provide a more aggressive discounting than in the original PDCFR+. Additionally, we set $\beta=1.5$. Other settings in APDCFR+ remain the same as in PDCFR+, such as the weighted averaging method in averaging strategies.
Finally, we introduce Stable APDCFR+, which updates the strategy at each infoset $I\in\mathcal{I}$ via
$\hat{\bm{R}}^{t}_{I}=\left[\frac{t^{\beta}}{\kappa+t^{\beta}}\bm{R}^{t}_{I}+\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\right]^{+},\ \bm{R}^{t+1}_{I}=\left[\frac{t^{\beta}}{\kappa+t^{\beta}}\bm{R}^{t}_{I}+\bm{r}^{t}_{I}\right]^{+},$ (42)
$\bm{R}^{t+1}_{I}=\begin{cases}\bm{1}&\text{if }\bm{R}^{t+1}_{I}\leq\bm{1},\\ \bm{R}^{t+1}_{I}&\text{else },\end{cases},\ \sigma^{t}_{i}(I)=\frac{\left[\hat{\bm{R}}^{t}_{I}\right]^{+}}{\|\left[\hat{\bm{R}}^{t}_{I}\right]^{+}\|_{1}}=\frac{\hat{\bm{R}}^{t}_{I}}{\|\hat{\bm{R}}^{t}_{I}\|_{1}},\ \alpha^{t}_{I}=\min\left(\sqrt{\frac{\sum_{\tau=1}^{t-1}\|\bm{r}^{\tau}_{I}-\bm{r}^{\tau-1}_{I}\|^{2}_{2}}{\sum_{\tau=1}^{t-1}\|\bm{R}^{\tau+1}_{I}-\bm{R}^{\tau}_{I}\|^{2}_{2}}},10\right),$
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, $\bm{r}^{0}_{I}=\bm{0}$, $\kappa=500$, and $\beta=1.5$.
Table 2: Number of Information Sets of Games.
Kuhn Poker Leduc Poker Battleship (3) Liar’s Dice (4) Liar’s Dice (5) Liar’s Dice (6)
12 936 81,027 1,024 5,120 24,576
Goofspiel (4) Goofspiel (5) Goofspiel (6) Subgame (3) Subgame (4)
162 2,124 34,482 69,184 43,240
@@ Appendix D Additional Experiments
Dynamics of $\alpha^{t}_{I}$ in APCFR+. We now present the dynamics of $\alpha^{t}_{I}$ in APCFR+. The experimental results are shown in Figures 6 and 7. Note that we output the average value of $\alpha^{t}_{I}$ across all $I\in\mathcal{I}$. Initially, $\alpha^{t}_{I}$ increases rapidly. After approximately 100 iterations, $\alpha^{t}_{I}$ stops increasing. A closer examination of the dynamics of $\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}$ and $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ is that this is due to the fact that the values of both quantities are significantly larger in the initial phase than at later stages (as observed by the results shown in below).
@@
Dynamics of $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}\text{ and }\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$. Experimental results are shown in Figures 8 and 9. We show the cumulative values across all infosets. We observe that $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is at least three times larger than $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$ for both PCFR+ and our algorithms. Intuitively, this phenomenon can be explained by the fact that the upper bound of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ is only a quarter of the upper bound of $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, which is also used to derive SAPCFR+ (as detailed in Section 4.2). Therefore, there must exists a sequence of $\alpha^{t}_{I}$ such that $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2},$ where the term in the right-hand side is the regret bound for PCFR+. Specifically, let $\alpha^{t}_{I}=\alpha>0$ be a positive constant. To satisfy the inequality $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha}+\alpha\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2},$ it suffices to ensure that $\sum_{t=1}^{T}\alpha\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\frac{\alpha\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha}$ (which comes from $\sum_{t=1}^{T}\alpha\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}-\sum_{t=1}^{T}\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha}$ ), or equivalently, $(1+\alpha)\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$. Given that $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is at least three times larger than $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$, we can choose $0<\alpha\leq 2$ to ensure $(1+\alpha)\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ holds.

### v2

@@ Appendix B An Alternative Regret Upper Bound of APCFR+
$R^{T}(I)\leq\sqrt{\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}}.$ (31)
@@
$R^{T}(I)\leq\min\left(\sqrt{{\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)}},\ \sqrt{\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}}\right).$ (32)
We employ Theorem 4.1 rather than Theorem B.1 in the main text because the regret upper bound presented in Theorem B.1 is typically significantly larger than that in Theorem 4.1. In our experiments (as shown in Figures 8, 9, 10, and 11), we observe that the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ consistently increases over time. In contrast, the value of $\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)$ tends to stabilize, exhibiting a flattening trend. We hypothesize that this phenomenon arises from the fact that even when ${\bm{r}}^{t}_{I}$ and $\bm{r}^{t-1}_{I}$ are close, the value of $\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ remains extremely large.
@@
$\leq$ $\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+$
$\quad\quad\quad\frac{1}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
@@
$\leq$ $\langle{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\frac{1}{1+\alpha^{t}_{I}}\langle\bm{r}^{t-1}_{I},\hat{\bm{\theta}}^{t}_{I}-{\bm{\theta}}^{t+1}_{I}\rangle+\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})\right)+$
$\quad\quad\quad\frac{1}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right).$
Combining Eq. (22), (24) and (35), we have
$R^{T}(I)\leq$ $\frac{1}{\eta}\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})+\sum^{T}_{t=1}\langle{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+$ (36)
$\quad\quad\quad\sum^{T}_{t=1}\frac{1}{1+\alpha^{t}_{I}}\langle\bm{r}^{t-1}_{I},\hat{\bm{\theta}}^{t}_{I}-{\bm{\theta}}^{t+1}_{I}\rangle+\sum^{T}_{t=1}\frac{1}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
@@
where the third inequality comes from that $\forall b,c,\rho>0,bc\leq b^{2}/(2\rho)+\rho c^{2}/2$ (in this case, $b=\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|_{2}$, $c=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}$, and $\rho=1/\eta$), and the last inequality is from $\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|^{2}_{2}/2$. Then, from Eq. (27) ($\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})\leq\frac{1}{2}$), we get
@@
Appendix C Details of APDCFR+
Next, we introduce APDCFR+. At each iteration $t$, the strategy is updated at each infoset $I\in\mathcal{I}$ as follows:
$\hat{\bm{R}}^{t}_{I}=\left[\frac{\lambda t^{\beta}}{\kappa+t^{\beta}}\bm{R}^{t}_{I}+\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\right]^{+},\ \bm{R}^{t+1}_{I}=\left[\bm{R}^{t}_{I}+\frac{\lambda t^{\beta}}{\kappa+t^{\beta}}\bm{r}^{t}_{I}\right]^{+},$ (40)
$\sigma^{t}_{i}(I)=\frac{\left[\hat{\bm{R}}^{t}_{I}\right]^{+}}{\|\left[\hat{\bm{R}}^{t}_{I}\right]^{+}\|_{1}}=\frac{\hat{\bm{R}}^{t}_{I}}{\|\hat{\bm{R}}^{t}_{I}\|_{1}},\ \alpha^{t}_{I}=\min\left(\sqrt{\frac{\sum_{\tau=1}^{t-1}\|\bm{r}^{\tau}_{I}-\bm{r}^{\tau-1}_{I}\|^{2}_{2}}{\sum_{\tau=1}^{t-1}\|\bm{R}^{\tau+1}_{I}-\bm{R}^{\tau}_{I}\|^{2}_{2}}},9\right),$
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, and $\bm{r}^{0}_{I}=\bm{0}$. In updating $\alpha^{t}_{I}$, we opt for $\alpha_{max}=9$ instead of $5$ as our observations suggest that $9$ yields superior performance within HUNL Subgames. Furthermore, we utilize the expression $\frac{\lambda t^{\beta}}{\kappa+t^{\beta}}$ for discounting accumulated regrets, in contrast to the $\frac{t^{\beta}}{1+t^{\beta}}$ used in DCFR. By setting different values for $\lambda$ and $\kappa$, we achieve a more aggressive discounting strategy than that in the original DCFR. Specifically, within this study, we assign $\lambda=20$ and $\kappa=500$. This configuration implies that for $\bm{R}^{t+1}_{I}$, the impact of $\bm{r}^{t}_{I}$ near the initial stages—where $t$ approaches 11—is relatively extremely smaller compared to employing $\frac{t^{\beta}}{1+t^{\beta}}$. Formally, by setting $\lambda=20$ and $\kappa=500$, the range of the values of $\frac{\lambda t^{\beta}}{\kappa+t^{\beta}}$ is $\frac{20}{501},20]$, while the range of the values of $\frac{t^{\beta}}{1+t^{\beta}}$ is $\frac{1}{2},1]$. APDCFR+ uses the following weighted averaging strategy:
$\bar{\sigma}^{t}=\left(\frac{t-1}{t}\right)^{2.5}\bar{\sigma}^{t-1}+\sigma^{t}.$ (41)
Notably, the parameters $\alpha_{\max}=9$, $\lambda=20$, $\kappa=500$, and $\beta=1.5$ are selected from the ranges $[5,7,9]$, $[10,20,50,100]$, $[100,500,1000]$, and $[1.0,1.5,2.0,2.5]$, respectively. The parameter tuning of APDCFR+ is conducted exclusively on HUNL subgames to identify the optimal configuration.
@@ Appendix D Additional Experiments
Game #Histories #Infosets #Terminal histories #Depth #Max size of infosets
Kuhn Poker 58 12 30 6 2
Leduc Poker 9,457 936 5,520 12 5
Battleship (3,2,3) 732,607 81,027 552,132 9 7
Battleship (4,3,2) 5,462,407 58,159 4,966,176 7 17
Liar’s Dice (4) 8,181 1,024 4,080 12 4
Liar’s Dice (5) 51,181 5,120 25,575 14 5
Goofspiel (4) 1,077 162 576 7 14
Goofspiel (5) 26,931 2,124 14,400 9 46
Subgame 3 398,112,843 69,184 261,126,360 10 1,980
Subgame 4 244,005,483 43,240 158,388,120 8 1,980
Leduc Poker (5) 55,361 2,760 32,760 12 9
Leduc Poker (9) 371,809 9,288 221,544 12 17
Leduc Poker (13) 1,179,777 19,656 704,600 12 25
Table 3: Sizes of the games.
Sizes of the Games. Before introducing our additional experiments, we present the sizes of the games used in our experiments, as shown in Table 3. In this table, #Histories denotes the number of histories in the game tree, while #Infosets represents the number of infosets in the game tree. #Terminal histories indicates the number of terminal histories in the game tree. #Depth defines the depth of the game tree, which is the maximum number of actions in a single history. Lastly, #Max size of infosets indicates the maximum number of histories that belong to the same infoset.
Game PCFR+ APCFR+ SAPCFR+
kuhn Poker 0.0020 0.0033 0.0021
Leduc Poker 0.4252 0.6423 0.4408
Battleship (3,2,3) 50.1909 58.4390 50.8602
Battleship (4,3,2) 103.2314 116.2511 102.4161
Liar’s Dice (4) 0.4327 0.5690 0.4308
Liar’s Dice (5) 3.8178 4.0365 3.7922
Goofspiel (4) 0.0290 0.0393 0.0279
Goofspiel (5) 0.9323 1.9760 0.9430
Subgame 3 278.7428 297.3784 276.9873
Subgame 4 245.1753 254.4792 249.0042
Table 4: Comparison of running times (in minutes).
Running times. We now compare the running time of our algorithms with that of PCFR+, both executed with the same number of iterations (i.e., 5000). The experimental results are shown in Table 4. We observe that the running time of APCFR+ is slightly higher compared to PCFR+, primarily due to the additional $\alpha^{t}_{I}$ learning process in APCFR+, which increases its running time. However, we also note that the running time of SAPCFR+ is nearly identical to that of PCFR+, as the only difference between their implementations is a single line of code, which does not alter the computational complexity. Importantly, the computational complexity remains exactly the same, even with no change in the constant factors.
@@
Dynamics of $\alpha^{t}_{I}$ in APCFR+. We now present the dynamics of $\alpha^{t}_{I}$ in APCFR+. The experimental results are shown in Figures 4 and 5. Note that we output the average value of $\alpha^{t}_{I}$ across all $I\in\mathcal{I}$. Initially, $\alpha^{t}_{I}$ increases rapidly. After approximately 100 iterations, $\alpha^{t}_{I}$ stops increasing. A closer examination is that the values of $\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}$ and $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ are significantly larger in the initial phase than at later stages (as observed by the results shown in below).
@@
Dynamics of $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}\text{ and }\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$. Experimental results are shown in Figures 6 and 7. We show the cumulative values across all infosets. We observe that $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is at least three times larger than $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$ for both PCFR+ and our algorithms. Intuitively, this phenomenon can be explained by the fact that the upper bound of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ is only a quarter of the upper bound of $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, which is also used to derive SAPCFR+ (as detailed in Section 4.2). Therefore, there must exists a sequence of $\alpha^{t}_{I}$ such that $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2},$ where the term in the right-hand side is the regret bound for PCFR+. Specifically, let $\alpha^{t}_{I}=\alpha>0$ be a positive constant. To satisfy the inequality $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha}+\alpha\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2},$ it suffices to ensure that $\sum_{t=1}^{T}\alpha\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\frac{\alpha\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha}$ (which comes from $\sum_{t=1}^{T}\alpha\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}-\sum_{t=1}^{T}\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha}$ ), or equivalently, $(1+\alpha)\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$. Given that $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is at least three times larger than $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$, we can choose $0<\alpha\leq 2$ to ensure $(1+\alpha)\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\leq\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ holds.